a and b are two stations 390 km apart.a train starts from a at 10 am and travels towards b at 65kmph.another train starts from b at 11 am .and travel towards a at 35kmph.at what time do they meet?
Answer by stanbon(75887)

a and b are two stations 390 km apart.a train starts from a at 10 am and travels towards b at 65kmph.another train starts from b at 11 am .and travel towards a at 35kmph.at what time do they meet?
------
a to b DATA:
rate = 65 kph ; time = x hrs ; distance = 65x km
-------
b to a DATA:
rate = 35 kph ; time = x-1 hrs ; distance = 35(x-1) km
---------
Equation:
65x + 35x-35 = 390 km
------------
100x = 355 km
----
x = 3.55 hrs = 3hrs 33 minutes
------
Time they meet: 10:00 AM +3 hrs 33 minutes = 1:33 PM
===========================
Cheers,
Stan H.
==============